OpenStudy (anonymous):

Ok so in order to solve this problem #54 you have to use the result from problem #53. #53. cos 2a = cos^2a - sin^2a HINT: WRITE TAN 2A AS TAN (A+A)........................#54. tan (pi/4 +a) - tan (pi/4 -a) = 2tan2a

OpenStudy (anonymous):

@perl

OpenStudy (er.mohd.amir):

tan(pi/4+a)={tan pi/4+tan a}/{1-tan pi/4*tan a} and tan pi/4=1 tan(pi/4-a)={tan pi/4-tan a}/{1+tan pi/4*tan a}

OpenStudy (anonymous):

I don't understand

OpenStudy (mathmale):

Loving: Your question needs clarification. Exactly what are you asking? looking for? Please remember that others don't have access to your Problems #53, #54, etc., and may need to have that access to understand what your goal or goals are.

OpenStudy (anonymous):

I gave question #53 and #54

OpenStudy (anonymous):

Looking for exact values

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@DaBest21

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

um i dont know srry

OpenStudy (littlebird):

Are you trying to prove that one side is equal to the other?

OpenStudy (triciaal):

can you put question 54 here by itself

OpenStudy (anonymous):

#54. tan (pi/4 +a) - tan (pi/4 -a) = 2tan2a

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

now what are the directions? does it say prove that.....

OpenStudy (anonymous):

hold on brb

OpenStudy (anonymous):

I think we have to solve 53 first

OpenStudy (anonymous):

It wants you to verify each identity

OpenStudy (triciaal):